# ComFairGNN: Community Fair Graph Neural Network

### v1

@@ Abstract.
Graph Neural Networks (GNNs) have become the leading approach for addressing graph analytical problems in various real-world scenarios. However, GNNs may produce biased predictions against certain demographic subgroups due to node attributes and neighbors surrounding a node. Most current research on GNN fairness focuses predominantly on debiasing GNNs using oversimplified fairness evaluation metrics, which can give a misleading impression of fairness. Understanding the potential evaluation paradoxes due to the complicated nature of the graph structure is crucial for developing effective GNN debiasing mechanisms. In this paper, we examine the effectiveness of current GNN debiasing methods in terms of unfairness evaluation. Specifically, we introduce a community-level strategy to measure bias in GNNs and evaluate debiasing methods at this level. Further, We introduce ComFairGNN, a novel framework designed to mitigate community-level bias in GNNs. Our approach employs a learnable coreset-based debiasing function that addresses bias arising from diverse local neighborhood distributions during GNNs neighborhood aggregation. Comprehensive evaluations on three benchmark datasets demonstrate our model’s effectiveness in both accuracy and fairness metrics.
paradox, measuring group fairness, evaluation metrics, graph neural network, bias mitigation
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
1. Introduction
@@
In today’s interconnected world, graph learning supports various real-world applications, such as social networks, recommender systems , and knowledge graphs (Bourigault et al., 2014; Nickel et al., 2015). Graph Neural Networks (GNNs) are powerful for graph representation learning and are used in tasks like node classification and link prediction by aggregating neighboring node information. However, GNNs often overlook fairness, leading to biased decisions due to structural bias and attribute bias like gender, race, and political ideology. These bias can cause ethical dilemmas in critical contexts, such as job candidate evaluations, where a candidate might be favored due to shared ethnic background or mutual acquaintances.
@@
Figure 1. An illustration of polarized graph: Two nodes $a$ and $b$ from different communities of the graph share identical labels; GNN learns different embedding for $a$ and $b$ due to structural bias.
@@
To tackle the outlined issue, multiple methods were suggested to evaluate and mitigate the fairness of node representation learning on graphs. Most of these methods aim to learn node representations that can yield statistically fair predictions across subgroups defined based on sensitive attributes (Lambrecht and Tucker, 2019; Kamishima and Akaho, 2017; Saxena et al., 2022). Choosing the right metric to evaluate bias in graphs inherently depends on the specific task at hand. The existing methods employed graph-level metrics originally designed for other purposes to address fairness concerns. The primary difficulty in addressing fairness within graphs arises from the frequent correlation between the graph’s topology and the sensitive attribute we aim to disregard. However, due to the complicated nature of graph structures, conducting graph-level fairness evaluations and comparisons of different methods is not as straightforward as commonly reported in the existing methods.
Recently works (Xu et al., 2018; Kusner et al., 2017; Gölz et al., 2019; von Kügelgen et al., 2021) point out that the reported scores in prior works might not reflect the true performance in the real-world application because of various evaluation issues and the nature of the dataset. For example, Simpson’s paradox (Blyth, 1972), highlights the context of the graduate admissions bias at UC Berkeley in 1973. Historically, it was believed that 44% of the 8,442 male applicants were admitted, compared to 35% of the 4,342 female applicants, indicating a gender bias. However, a closer look reveals that four out of the six departments had a higher acceptance rate for women, while two departments had a higher acceptance rate for men. A recent example of Simpson’s paradox during the COVID-19 era illustrates its implications in health policy. In 2020, initial data showed a higher COVID-19 case fatality rate in Italy than in China. However, this analysis was confounded by age distribution differences. When analyzed by age groups, the case fatality rate was actually higher in China than in Italy within each age group (von Kügelgen et al., 2021). The aforementioned works are not focused on fairness for machine learning on graphs.
The majority of real-world graphs are polarized into communities due to some sensitive features such as age, gender, race, ideology, and socio-economic status, as shown in figure 1. A community is a group of nodes that are highly interconnected within the group and have fewer connections to nodes outside the group. For instance, research indicates that the dynamics of interaction among communities within United States Twitter networks are significantly shaped by their political allegiances (Guerra et al., 2013). Within the country, the flow of communication between communities in the Northeast and West Coast is largely driven by Democratic supporters, contrasting with the influence of Republican supporters. Conversely, in Texas, Florida, and Iowa, Republican supporters tend to dominate the network flow. This suggests a geographical alternation bubble in the dominance of network flows between Democrat and Republican communities. Consider a scenario where social media users are categorized into those who like basketball and those who like football. The demographics of those subgroups may vary across different regions despite the existence of some individuals who participate in both sports.
@@
In this paper, we investigate the important problem of community fairness in graph neural networks in the context of node classification. On the one hand, simply employing node attribute and their structure debiasing function may result in misleading fairness evaluation and loss of information (Wu et al., 2019). Communities are a manifestation of the local neighborhood structures, which are fundamentally different from a sensitive node attribute (Rahman et al., 2019). Particularly, in GNNs, each node learns by aggregation of information from the neighboring nodes, thus identical label nodes with varying neighborhood distribution as illustrated in figure  1, giving rise to structural bias in the learning process. Moreover, the bias may be further amplified by multi-layered recursive neighborhood aggregation in GNNs. Although current debiasing methods primarily address attribute-based node fairness, they fail to tackle the fundamental bias inherent in GNN neighborhood aggregation. Consequently, these approaches are unable to effectively mitigate local community-level biases arising from disparate neighborhood structures.
@@
To address community fairness, we introduce the concept of local structural fairness at the community level for all nodes in the graph. This approach recognizes that the broader structural bias affecting nodes within a community emerges from the GNN’s aggregation process, which is influenced by the diverse distribution of local neighborhoods. We present ComFairGNN, a novel community-fair Graph Neural Network compatible with any neighborhood aggregation-based GNN architecture. Specifically, the coreset-based debiasing function targets the neighborhood aggregation of GNN operation, aiming to bring identical label nodes closer in the embedding space. This approach mitigates structural disparities across communities, leading to fairer representations for nodes with identical labels.
@@
2. PRELIMINARIES
In this section, we introduce the structural-based clustering setting to assess our research question: to what extent are the existing graph fairness evaluation metrics prevalent in fair decisions at the community level and prevent the corresponding types of fairness paradoxes?
@@ 2. PRELIMINARIES
Let’s first introduce the notations. A graph $G=(V,E)$ consists of $n$ nodes $V$ and edges $E\subseteq V\times V$. The node-wise feature matrix $X\in\mathbb{R}^{n\times k}$ has $k$ dimensions for raw node features, with each row $\mathbf{x}_{i}$ representing the feature vector for the $i$-th node. The binary adjacency matrix is $A\in\{0,1\}^{n\times n}$, and the learned node representations are captured in $H\in\mathbb{R}^{n\times d}$, where $d$ is the latent dimension size and $\mathbf{h}_{i}$ is the representation for the $i$-th node. The sensitive attribute $s\in\{0,1\}^{n}$ classifies nodes into demographic groups, with edges $e_{uv}$ being intragroup if $s_{u}=s_{v}$ and intergroup otherwise. The function $N(u)$ returns the set of neighbors for a node $u$, providing a structural node embedding. Clusters based on this embedding are denoted as $C$, with $C_{i}\subseteq V$ for nodes in the $i$-th cluster and $c$ as the cluster count. Each $C_{i}$ represents a community and $f$ measures the statistical notion of fairness across different communities in the graph.
2.1. Graph Neural Network
@@
Graph Neural Networks (GNNs) utilize a layered architecture for neighborhood aggregation, where nodes progressively integrate information from their immediate neighbors across multiple iterations. For the $l$-th layer, the representation of a node $v$, denoted as $\mathbf{h}_{v}^{l}\in\mathbb{R}^{d_{l}}$, is formulated as follows:
@@
(1) $\mathbf{h}_{v}^{l}=\sigma\left(\text{AGGR}\left(\{\mathbf{h}_{u}^{l-1}\mid u\in\mathcal{N}_{v}\};\omega^{l}\right)\right)$
@@
where $d_{l}$ is the dimension of node representations in the $l$-th layer, $\text{AGGR}(\cdot)$ denotes an aggregation function such as mean-pooling (Kipf and Welling, 2016) or self-attention (Veličković et al., 2017), $\sigma$ is an activation function, $\mathcal{N}_{v}$ denotes the set of neighbors of $v$, and $\omega^{l}$ denotes the learnable parameters in layer $l$. The initial layer (layer 0) utilizes the raw input features as node representations, expressed mathematically as $\mathbf{h}_{v}^{0}=\mathbf{x}_{v}$, where $\mathbf{x}_{v}$ denotes the input attribute vector for node $v$.
2.2. Node Clustering on Structural Embedding
By integrating node2vec (Grover and Leskovec, 2016) embeddings with $k$-means clustering (MacQueen et al., 1967), we can effectively partition the graph into communities based on the structural similarities captured by the embeddings.
Graph embedding using node2vec is a mapping that assigns each node in a graph to a vector in a $d$-dimensional space, denoted as $f:U\rightarrow\mathbb{R}^{d}$, where $d$ is a hyperparameter representing the number of dimensions in the vector space. For each node $u\in U$ in a graph $G$, multiple random walks are performed, as specified by the hyperparameter walk_num. This process generates a list of sequences, each comprising node IDs from individual random walks. A single random walk follows the standard approach in unweighted graphs, where at each step, the next node is selected uniformly at random from the current node’s neighbors. The hyperparameters walk_num and walk_len control the number of walks and the length of each walk, respectively. Next, node2vec uses the generated sequences to train a neural network and learn embedding vectors. Let us define a network neighborhood $N_{s}(u)$ as the set of nodes preceding and succeeding node $u$ in the generated walk sequences. The objective function is defined as:
(2) $\arg\max_{f}\prod_{u\in U}\prod_{u^{\prime}\in N_{s}(u)}P(u^{\prime}\mid f(u))$
$P(u^{\prime}\mid f(u))$ is the conditional probability represented as softmax.
Once the node embeddings are learned using node2vec, the next step involves clustering these embeddings to identify groups of similar nodes. Our primary focus for clustering is on $k$-means clustering, selected for its efficient linear runtime. $k$-means aims to minimize the total sum of squared errors, also known as inertia or within-cluster sum of squares (Shiao et al., 2023). This objective can be formally expressed as:
(3) $\arg\min\sum_{i=1}^{C}{\sum_{x\in C_{i}}\text{DIST}(x,\mu_{i})}.$
2.3. Group Fairness Metrics on Graphs
We specifically target group fairness scenarios where both the node label and the sensitive attribute are binary. Group fairness is the most prevalent and extensively studied form of fairness. In the context of group fairness, we consider a model or algorithm to be fair when its outputs exhibit no bias against any demographic group of sensitive attributes. In this work, we primarily focus on assessing the evaluation of existing graph fairness methods utilizing the two widely used metrics of group fairness, i.e., demographic parity (Dwork et al., 2012) and equal opportunity (Hardt et al., 2016).
Let $\hat{y}$ denote the prediction output of the classifier $f(\cdot)$. $y\in\{0,1\}$ and $s\in\{0,1\}$ represent the ground truth and the sensitive attribute, respectively.
Demographic Parity: Demographic parity, also referred to as statistical parity, necessitates that the prediction $\hat{y}$ is independent of the sensitive attribute $s$, symbolized as $\hat{y}\perp s$. Formally, demographic parity aims to have:
(4) $P(\hat{y}|s=0)=P(\hat{y}|s=1)$
Equal Opportunity: Equal opportunity involves an identical true positive rate across all demographic groups. This concept can be expressed as:
(5) $P(\hat{y}=1|y=1,s=0)=P(\hat{y}=1|y=1,s=1)$
During experiments, demographic parity and equal opportunity are utilized to assess fairness performance as below:
(6) $\Delta{\text{SP}}=\left|P(\hat{y}=1|s=0)-P(\hat{y}=1|s=1)\right|$
(7) $\Delta{\text{EO}}=\left|P(\hat{y}=1|y=1,s=0)-P(\hat{y}=1|y=1,s=1)\right|$
3. PROPOSED METHOD
The overall framework of the ComFairGNN is illustrated in Figure 2. We first design a strategy of communities using a clustering algorithm based on structural contrast in Figure 2(c) to evaluate the effectiveness of the existing fairness methods and our proposed method at the community level. To facilitate the debiasing function, we split nodes in each community into two subgroups based on their sensitive feature and sample coreset nodes from each community based on their neighborhood homophily distribution ratio Figure 2(c). The coreset nodes contrast each other based on their output similarity to mitigate structural bias at the community level Figure 2(d). More specifically, the sampled coreset based debiasing function aims to balance the neighborhood distribution of nodes in different communities and remove bias by improving the output similarity of the nodes that have the same label. Thus, by trying to maximize the similarity of nodes in the coreset with the same label, the GNN is able to debias the neighborhood aggregation and balance the two groups. The GNN is optimized by the task loss and the coreset fairness loss Figure 2(e).
Figure 2. Overall framework of ComFairGNN
3.1. Community level Structural Contrast
To debias nodes in a community with varying label neighborhood distribution, we propose a debiasing function based on the coreset sample nodes $c$, which are selected from different communities. Coreset selection methods aim to identify crucial subsets of training examples by employing specific heuristic criteria (Toneva et al., 2018). These subsets are considered essential for effectively training the model. Our study proposes a new method of graph coreset selection method to debias GNNs, addressing the local structural bias in their aggregation process. The coreset nodes are selected equally from each community for both groups to learn the debiasing function by contrasting between two groups of nodes, namely, subgroups $\mathcal{S}_{0}$ and $\mathcal{S}_{1}$. Note that GNNs neighborhood aggregation in Eq. 1, can only access its one-hop local contexts in each layer. The majority of real-world networks of nodes in a community tend to be from the same subgroup influenced by the sensitive attribute. Thus, it is a natural choice to select the coreset sample nodes of both groups from each community to balance the structural imbalance between different communities.
(8) $c=\bigcup_{k=1}^{K}C_{k_{S_{0}}}\cup\bigcup_{k=1}^{K}C_{k_{S_{1}}}\subseteq V$
@@ 3.1. Community level Structural Contrast
To minimize the structural bias between the two groups in different communities, we contrast the output similarity of nodes with the same label in the coreset. This tends to maximize the similarity of nodes of the same label in different communities. This strategy enables GNNs to implicitly eliminate the structural bias between different communities.
3.2. Coreset Selection for GNNs Debiasing
To fundamentally eliminate the structural bias of GNNs neighborhood aggregation, which is the key operation in GNNs, we propose a neighborhood homophily distribution ratio to select the sample coreset from different communities.
Node neighborhood homophily distribution ratio: The homophily ratio in graphs is typically defined based on the similarity between connected node pairs. In this context, two nodes are considered similar if they share the same node label. The formal definition of the homophily ratio is derived from this intuition, as follows (Ma et al., 2021).
Definition 0 (Node Homophily Ratio).
@@ 3.2. Coreset Selection for GNNs Debiasing
(9) $h(G,\{y_{i};i\in\mathcal{V}\})=\frac{1}{|n|}\sum_{(i,k)\in\mathcal{E}}f(y_{i}=y_{k}),$
@@
As shown in figure 3, the neighborhood distribution ratio for identical label nodes in pokec-z dataset varies largely. If we consider that all nodes that share the same label have similar initial node features, then cleary, after GNN 1-hop neighborhood aggregation, the identical label nodes will have different output embedding. This inability to differentiate the embedding space of identical nodes in different communities of the graph can potentially lead the GNN to be biased towards certain subgroups. Considering this, the debiasing coreset nodes are sampled based on their neighborhood distribution from all communities of the graph.
Figure 3. Edge neighborhood distribution ratio for Pokec-z dataset for nodes with Label 0 and Label 1: GNN learns biased embedding space for different neighborhood distribution of identical nodes in different communities.
Fairness-aware Graph Coreset: aims to identify a representative subset of training nodes $V_{w}\subset V_{\text{train}}=[n_{t}]$, where $|V_{w}|\leq c\ll n_{t}$, along with an associated set of sample weights. This selection is designed to approximate the training loss $\mathcal{L}(\theta)$ across all $\theta\in\Theta$, while maintaining fairness considerations. The graph coreset selection problem is formulated as:
(10) $\min_{w\in\mathcal{W}}\max_{\theta\in\Theta}\left|\sum_{i\in[n_{t}]}w_{i}\cdot\ell([f_{\theta}(A,X)]_{i},y_{i})-\mathcal{L}(\theta)\right|$
which minimizes the worst-case error for all $\theta$. Both the graph adjacency matrix $A$ and node feature matrix $X$ are still needed to compute the full-graph embeddings $f_{\theta}(A,X)$ in the coreset loss. Since the goal of the fairness-aware graph, coreset is to debias the neighborhood aggregation of GNNs at the community level, we consider the coreset samples of both groups of $S_{0}$ and $S_{1}$ from all communities of the graph.
3.3. Training Objective and Constraints:

### v2

@@ Abstract
Graph Neural Networks (GNNs) have become the leading approach for addressing graph analytical problems in various real-world scenarios. However, GNNs may produce biased predictions against certain demographic subgroups due to node attributes and neighbors surrounding a node. Understanding the potential evaluation paradoxes due to the complicated nature of the graph structure is crucial for developing effective GNN debiasing mechanisms. In this paper, we examine the effectiveness of current GNN debiasing methods in terms of unfairness evaluation. Specifically, we introduce a community-level strategy to measure bias in GNNs and evaluate debiasing methods at this level. Further, We introduce ComFairGNN, a novel framework designed to mitigate community-level bias in GNNs. Our approach employs a learnable coreset-based debiasing function that addresses bias arising from diverse local neighborhood distributions during GNNs neighborhood aggregation. Comprehensive evaluations on three benchmark datasets demonstrate our model’s effectiveness in both accuracy and fairness metrics.
Keywords:
fairness graph neural network bias evaluation metrics
1 Introduction
@@
In today’s interconnected world, graph learning supports various real-world applications, such as social networks, recommender systems , and knowledge graphs [3, 16]. Graph Neural Networks (GNNs) are powerful for graph representation learning and are used in tasks like node classification and link prediction by aggregating neighboring node information. However, GNNs often overlook fairness, leading to biased decisions due to structural bias and attribute bias like gender, race, and political ideology. These bias can cause ethical dilemmas in critical contexts, such as job candidate evaluations, where a candidate might be favored due to shared ethnic background or mutual acquaintances.
@@
Figure 1: Illustration of graph polarization: Nodes $a$ and $b$ from different communities share identical labels (same colors), but GNNs learn different embeddings due to varying neighborhood label distributions.
@@
To tackle the outlined issue, multiple methods were suggested to evaluate and mitigate the fairness of node representation learning on graphs. Most of these methods aim to learn node representations that can yield statistically fair predictions across subgroups defined based on sensitive attributes. Choosing the right metric to evaluate bias in graphs inherently depends on the specific task at hand. The existing methods employed graph-level metrics originally designed for other purposes to address fairness concerns. The primary difficulty in addressing fairness within graphs arises from the frequent correlation between the graph’s topology and the sensitive attribute we aim to disregard. However, due to the complicated nature of graph structures, conducting graph-level fairness evaluations and comparisons of different methods is not as straightforward as commonly reported in the existing methods.
Real-world graphs often polarize into communities based on sensitive attributes (e.g., age, gender, race, ideology), as shown in Figure 1, where communities form through dense internal connections and sparse external links. For example, U.S. Twitter networks show distinct political segregation [9], where Democratic supporters dominate Northeast and West Coast interactions while Republican supporters influence Texas, Florida, and Iowa networks. Similarly, sports preferences on social media can reflect regional demographic differences.
@@
In this paper, we investigate the critical issue of community fairness in graph neural networks (GNNs) within the context of node classification. Simply using node attributes and structural information as a debiasing mechanism can lead to misleading fairness evaluations and loss of important information [23]. Communities, which reflect local neighborhood structures, often stem from sensitive node attributes [17]. Specifically, GNNs generate node representations by aggregating information from neighboring nodes, which can result in nodes with identical labels but differing neighborhood label distributions, as illustrated in Figure 1. This leads to structural biases in the learning process. Moreover, such biases may be further amplified through multi-layered recursive neighborhood aggregation in GNNs. While current debiasing methods primarily focus on attribute-based fairness, they fail to address the underlying bias introduced by neighborhood aggregation. As a result, these methods are ineffective at mitigating community-level biases that arise from disparate neighborhood structures.
@@
To address community fairness, we introduce the concept of local structural fairness at the community level for all nodes in the graph. This approach recognizes that the broader structural bias affecting nodes within a community emerges from the GNN’s aggregation process, which is influenced by the diverse distribution of local neighborhoods. We present ComFairGNN, a novel community-fair Graph Neural Network compatible with any neighborhood aggregation based GNN architecture. Specifically, the coreset-based debiasing function targets the neighborhood aggregation of GNN operation, aiming to bring identical label nodes in different communities closer in the embedding space. This approach mitigates structural disparities across communities, leading to fairer representations for nodes with identical labels.
@@
2 Preliminaries
@@
We first introduce the notations. A graph $G=(V,E)$ consists of $n$ nodes and edges $E\subseteq V\times V$. The node-wise feature matrix $X\in\mathbb{R}^{n\times k}$ has $k$ dimensions for raw node features, with each row $\mathbf{x}_{i}$ representing the feature vector for the $i$-th node. The binary adjacency matrix is $A\in\{0,1\}^{n\times n}$, and the learned node representations are captured in $H\in\mathbb{R}^{n\times d}$, where $d$ is the latent dimension size and $\mathbf{h}_{i}$ is the representation for the $i$-th node. The sensitive attribute $s\in\{0,1\}^{n}$ classifies nodes into demographic groups, with edges $e_{uv}$ being intragroup if $s_{u}=s_{v}$ and intergroup otherwise. The function $N(u)$ returns the set of neighbors for a node $u$, providing a structural node embedding. Clusters based on this embedding are denoted as $C$, with $C_{i}\subseteq V$ for nodes in the $i$-th cluster and $c$ as the cluster count. Each $C_{i}$ represents a community and $f$ measures the statistical notion of fairness across different communities in the graph.
2.1 Graph Neural Network
@@
Graph Neural Networks (GNNs) utilize neighborhood aggregation across multiple layers. At layer $l$, the representation of node $v$, denoted as $\mathbf{h}_{v}^{l}\in\mathbb{R}^{d_{l}}$, is:
@@
$\mathbf{h}_{v}^{l}=\sigma\left(\text{AGGR}\left(\{\mathbf{h}_{u}^{l-1}\mid u\in\mathcal{N}_{v}\};\omega^{l}\right)\right)$ (1)
@@
where $\text{AGGR}(\cdot)$ is an aggregation function like mean-pooling [12] or self-attention [21], $\sigma$ is an activation function, $\mathcal{N}_{v}$ is $v$’s neighborhood set, and $\omega^{l}$ contains learnable parameters. The initial representation is $\mathbf{h}_{v}^{0}=\mathbf{x}_{v}$, where $\mathbf{x}_{v}$ is the input feature vector.
2.2 Node Clustering on Structural Embedding
We partition the graph into communities using node2vec [8] embeddings combined with $k$-means clustering [15]. Node2vec maps each node $u\in U$ to a $d$-dimensional vector space $f:U\rightarrow\mathbb{R}^{d}$. For each node, multiple random walks are performed to capture the network neighborhood $N_{s}(u)$, maximizing the objective:
$\arg\max_{f}\prod_{u\in U}\prod_{u^{\prime}\in N_{s}(u)}P(u^{\prime}\mid f(u))$ (2)
where $P(u^{\prime}\mid f(u))$ is computed using softmax. These structural embeddings are then clustered using $k$-means, chosen for its computational efficiency. The $k$-means algorithm minimizes the within-cluster sum of squares [18]:
$\arg\min\sum_{i=1}^{C}{\sum_{x\in C_{i}}\text{DIST}(x,\mu_{i})}$ (3)
where $C_{i}$ represents the $i$-th cluster and $\mu_{i}$ its centroid.
3 Proposed Method
Figure 2 illustrates the ComFairGNN framework. We first identify communities using structural-based clustering (Fig. 2(c)). For debiasing, we divide each community’s nodes into two subgroups based on sensitive features and sample coreset nodes according to neighborhood homophily ratios. These coreset nodes contrast against each other to mitigate structural bias (Fig. 2(d)). The GNN optimizes both task performance and coreset fairness (Fig. 2(e)), debiasing neighborhood aggregation by maximizing similarity between same-labeled nodes.
Figure 2: Overall framework of ComFairGNN
3.1 Community Level Structural Contrast
We propose a debiasing function using coreset nodes $c$ selected from different communities. Following [20], coresets identify crucial training examples using heuristic criteria. We select nodes equally from subgroups $\mathcal{S}_{0}$ and $\mathcal{S}_{1}$ across communities. Since GNNs in Eq. 1 only access one-hop contexts and communities tend to be homogeneous in sensitive attributes, selecting coreset nodes from both groups per community helps balance structural disparities.
$C=\bigcup_{k=1}^{K}C_{k_{S_{0}}}\cup\bigcup_{k=1}^{K}C_{k_{S_{1}}}\subseteq V$ (4)
@@ 3.1 Community Level Structural Contrast
To minimize structural bias between groups across communities, we maximize the similarity of similarly-labeled nodes in the coreset, allowing GNNs to implicitly reduce community-based structural differences.
3.2 Coreset Nodes Selection for GNNs Debiasing
To fundamentally eliminate the structural bias of GNNs neighborhood aggregation, which is the key operation in GNNs, we propose to use the neighborhood homophily distribution ratio to select the sample coreset nodes from different communities.
Node neighborhood homophily distribution ratio: The homophily ratio in graphs is typically defined based on the similarity between connected node pairs. In this context, two nodes are considered similar if they share the same node label. The formal definition of the homophily ratio is derived from this intuition, as follows [14].
Definition 1 (Node Homophily Ratio)
@@ 3.2 Coreset Nodes Selection for GNNs Debiasing
$h(G,\{y_{i};i\in\mathcal{V}\})=\frac{1}{|n|}\sum_{(i,k)\in\mathcal{E}}f(y_{i}=y_{k}),$ (5)
@@
Figure 3 shows significant variation in neighborhood distribution ratios among same-labeled nodes in pokec-z. Since nodes with identical labels typically share similar features, GNN’s 1-hop aggregation produces different embeddings for these nodes. This embedding disparity across communities may bias the GNN. Therefore, we sample debiasing coreset nodes from all communities based on their neighborhood distributions.
Figure 3: Edge distribution bias in Pokec-z: GNN embeddings vary for identical nodes based on neighborhood labels
Fairness-aware Graph Coreset: We aim to identify a representative subset of training nodes $V_{w}\subset V_{\text{train}}=[n_{t}]$, where $|V_{w}|\leq c\ll n_{t}$, along with an associated set of sample weights. In nearest neighbor message passing models (such as GCN, GAT, and GIN), node $i$ only interacts with nodes $j$ that are within a shortest-path distance of $D$, denoted as $\{j\in[n]\mid d(i,j)<D\}$. Here, we assume node $i$ neighborhood distribution ratio section 3.2, to select the coreset nodes. This selection is designed to approximate the training loss $\mathcal{L}(\theta)$ across all $\theta\in\Theta$, while maintaining fairness considerations. The graph coreset selection problem is formulated as:
$\min_{w\in\mathcal{W}}\max_{\theta\in\Theta}\left|\sum_{i\in[n_{t}]}w_{i}\cdot\ell([f_{\theta}(A,X)]_{i},y_{i})-\mathcal{L}(\theta)\right|$ (6)
The coreset minimizes the maximum error across all $\theta$, requiring both $A$ and $X$ matrices to compute $f_{\theta}(A,X)$. As illustrated in algorithm 1, to debias GNN neighborhood aggregation at the community level, we sample from both groups $S_{0}$ and $S_{1}$ across all communities.
Algorithm 1 Fair Graph Coreset Selection
0:  $G=(V,E)$, features $X$, groups $S_{0}$, $S_{1}$, size $K$
0:  Coreset $C\subseteq V$Selected nodes}
1:  Initialize $C=\emptyset$, $H=\text{ENC}(G)$Initial embedding}
2:  for each community $k$ in $G$ do
3:     $n_{k}=\lfloor K\cdot|V_{k}|/|V|\rfloor$Community budget}
4:     $C_{0k}=\text{Sample}(V_{k}\cap S_{0},\lfloor n_{k}/2\rfloor)$
5:     $C_{1k}=\text{Sample}(V_{k}\cap S_{1},\lfloor n_{k}/2\rfloor)$
6:     $C=C\cup C_{0k}\cup C_{1k}$
7:  end for
8:  $Z=\text{PRD}(H)$
9:  Optimize $C$ to minimize Task_Loss + Fairness_Loss
10:  return  $C$
3.3 Training Objective and Constraints: